# Gerry Windiarto Mohamad Dunda Electronic and Computer Engineering

### v4

@@ Appendix A Proof of Theorem 1
$\leq(1-\rho\eta_{\lambda,t})\|\lambda_{t}-\hat{\lambda}_{S}(w_{t})\|^{2}+\eta_{\lambda,t}^{2}\|\nabla_{\lambda}F_{S}(w^{(t,E)})\|^{2}+\eta_{\lambda,t}^{2}B_{\lambda}^{2}$
$+2\eta_{\lambda,t}\sum_{i=1}^{N}p_{i}(F_{i,S}(w_{i}^{(t,E)},\hat{\lambda}_{S}(w_{t}))-F_{i,S}(w_{i}^{(t,E)},\lambda_{t}))$
$\leq(1-\rho\eta_{\lambda,t})\|\lambda_{t}-\hat{\lambda}_{S}(w_{t})\|^{2}-\frac{\eta_{\lambda,t}}{L}\|\nabla_{\lambda}F_{S}(w_{i}^{(t,E)},\lambda_{t})\|^{2}$
$+\eta_{\lambda,t}^{2}\|\nabla_{\lambda}F_{S}(w^{(t,E)})\|^{2}+\eta_{\lambda,t}^{2}B_{\lambda}^{2}$
$\leq(1-\rho\eta_{\lambda,t})\|\lambda_{t}-\hat{\lambda}_{S}(w_{t})\|^{2}+\eta_{\lambda,t}^{2}B_{\lambda}^{2}.$ (26)
The fourth inequality is from the $\rho$-strongly concave property of $F_{i,S}(w,\cdot)$. The last inequality is obtained by $\eta_{\lambda,t}\leq\frac{1}{L}$ By substituting (25) and (26) into (24), we get the final result. ∎
@@
Define $a_{t}=\mathbb{E}_{t}[R_{S}(w_{t})-R^{*}_{S}]$ and $b_{t}=\mathbb{E}_{t}\|\lambda_{t}-\hat{\lambda}_{S}(w_{t})\|^{2}$. For any non-increasing sequence $\{\nu_{t}>0\}$ and any $\epsilon>0$, we have the following relations,
$a_{t+1}+\nu_{t+1}b_{t+1}\leq k_{1,t}a_{t}+k_{2,t}\nu_{t}b_{t}+c_{t},$
@@
$k_{1,t}$ $=1-\frac{\mu\eta_{w,t}}{\beta^{2}}+\frac{L^{2}E\eta_{w,t}}{\mu}+\nu_{t}(1+\frac{1}{\epsilon})(\frac{4L^{2}\eta_{w,t}}{\rho^{2}}+\frac{4L^{4}E\eta_{w,t}^{2}}{\rho^{2}\mu})$
$k_{2,t}$ $=\frac{L^{2}\eta_{w,t}}{2\nu_{t}}+(1+\epsilon)(1-\rho\eta_{\lambda,t})$
$c_{t}$ $=\frac{\eta_{w,t}}{2}\zeta_{t}+\frac{\eta^{\prime}_{w,t}}{2}E^{2}B_{w}^{2}+\frac{\mu}{\beta^{2}}\eta^{\prime}_{w,t}\kappa^{2}$
$+\nu_{t}(1+\frac{1}{\epsilon})\frac{L^{2}\eta_{w,t}^{2}}{\rho^{2}}EB_{w}^{2}+\nu_{t}(1+\epsilon)\eta_{\lambda,t}^{2}B_{\lambda}^{2}$
$+(\frac{L^{2}E\eta_{w,t}}{\mu}-\frac{\mu}{\beta^{2}}\eta_{w,t}+\nu_{t}(1+\frac{1}{\epsilon})\frac{4L^{4}E\eta_{w,t}^{2}}{\rho^{2}\mu})\Gamma.$
@@
$k_{2,t}$ $=\frac{L^{2}\eta_{w,t}}{2\nu_{t}}+1-\frac{\rho\eta_{\lambda,t}}{2}.$
@@
$k_{1,t}$ $=1-\frac{\mu\eta_{w,t}}{\beta^{2}}+\frac{L^{2}E\eta_{w,t}}{\mu}+\frac{16L^{2}\eta_{w,t}^{2}}{\rho^{4}\eta_{\lambda,t}^{2}}(1+\frac{L^{2}E\eta_{w,t}}{\mu})$ (27)
@@
Let $\eta^{2}_{\lambda,t}\geq\frac{16L^{2}\eta_{w,t}}{\rho^{4}(\frac{\mu}{2\beta^{2}}-\frac{L^{2}E}{\mu})}(1+\frac{L^{2}E\eta_{w,t}}{\mu})$ and $\eta_{\lambda,t}\geq\frac{2\mu}{\beta^{2}\rho}\eta_{w,t}$. We get
$a_{t+1}+\nu_{t+1}b_{t+1}\leq(1-\frac{\mu}{2\beta^{2}}\eta_{w,t})(a_{t}+\nu_{t}b_{t})+\frac{\eta^{\prime}_{w,t}}{2}E^{2}B_{w}^{2}+\frac{\mu}{\beta^{2}}\eta^{\prime}_{w,t}\kappa^{2}$
$+\frac{4L^{2}\eta_{w,t}^{3}}{\rho^{3}\eta^{2}_{\lambda,t}}EB_{w}^{2}+\frac{L^{2}\eta_{\lambda,t}\eta_{w,t}B_{\lambda}^{2}}{\rho(1-\rho\eta_{\lambda,t})}$
$+(\frac{L^{2}E\eta_{w,t}}{\mu}-\frac{\mu}{\beta^{2}}\eta_{w,t}+\frac{16L^{6}E\eta_{w,t}^{3}}{\rho^{4}\mu\eta_{\lambda,t}^{2}})\Gamma.$ (29)
Choose $\eta_{w,t}=\frac{2\beta^{2}}{\mu t}$ and $\eta_{\lambda,t}=\mathcal{O}(\frac{1}{t^{2/3}})$ and multiply both sides of (29) by $t$ to get
$t(a_{t+1}+\nu_{t+1}b_{t+1})$ $\leq(t-1)(a_{t}+\nu_{t}b_{t})+\frac{2\kappa^{2}(L+\frac{L^{2}}{\rho})}{\beta^{2}t}$
$+\frac{E^{2}B_{w}^{2}(L+\frac{L^{2}}{\rho})}{\mu t}+(\Gamma+B_{w}^{2}+B_{\lambda}^{2})\mathcal{O}(\frac{1}{t^{2/3}})$
$+\mathcal{O}(\frac{1}{t^{4/3}})$ (30)
Applying the inequality inductively from $t=1$ to $T$ and divide both sides by $T$, we get
$\mathbb{E}[R_{S}(w_{T+1})-R^{*}_{S}]$ $\leq\frac{2\kappa^{2}(L+\frac{L^{2}}{\rho})}{\beta^{2}T^{2}}+\frac{E^{2}B_{w}^{2}(L+\frac{L^{2}}{\rho})}{\mu T^{2}}$
$+(\Gamma+B_{w}^{2}+B_{\lambda}^{2})\mathcal{O}(\frac{1}{T^{2/3}})+\mathcal{O}(\frac{1}{T^{4/3}}).$ (31)
Appendix B Additional Experiments on Different Levels of Heterogeneity

### v5

@@ Appendix A Proof of Theorem 1
$\leq\mathbb{E}\|\lambda_{t}-\lambda^{*}_{S}(w_{t})\|^{2}+2\eta_{\lambda,t}\langle\lambda_{t}-\lambda^{*}_{S}(w_{t}),\sum_{i=1}^{N}p_{i}\nabla_{\lambda}F_{i,S}(w^{(t,E)},\lambda_{t})\rangle$
$+\eta_{\lambda,t}^{2}\sum_{i=1}^{N}p_{i}\|\nabla_{\lambda}F_{S}(w_{i}^{(t,E)},\lambda_{t})\|^{2}+\eta_{\lambda,t}^{2}B_{\lambda}^{2}$
$\leq(1-\rho\eta_{\lambda,t})\mathbb{E}\|\lambda_{t}-\lambda^{*}_{S}(w_{t})\|^{2}$
$+2\eta_{\lambda,t}\sum_{i}p_{i}(F_{i,S}(w_{i}^{(t,E)},\lambda_{t})-F_{i,S}(w_{i}^{(t,E)},\lambda^{*}_{S}(w_{t})))$
$+\eta_{\lambda,t}^{2}\sum_{i=1}^{N}p_{i}\|\nabla_{\lambda}F_{S}(w_{i}^{(t,E)},\lambda_{t})\|^{2}+\eta_{\lambda,t}^{2}B_{\lambda}^{2}$
$\leq(1-\rho\eta_{\lambda,t})\mathbb{E}\|\lambda_{t}-\lambda^{*}_{S}(w_{t})\|^{2}-\frac{\eta_{\lambda,t}}{L}\sum_{i}p_{i}\|\nabla_{\lambda}F_{i,S}(w_{i}^{(t,E)},\lambda_{t})\|^{2}$
$+\eta_{\lambda,t}^{2}\sum_{i=1}^{N}p_{i}\|\nabla_{\lambda}F_{S}(w_{i}^{(t,E)},\lambda_{t})\|^{2}+\eta_{\lambda,t}^{2}B_{\lambda}^{2}$
$\leq(1-\rho\eta_{\lambda,t})\mathbb{E}\|\lambda_{t}-\lambda^{*}_{S}(w_{t})\|^{2}+\eta_{\lambda,t}^{2}B_{\lambda}^{2}.$ (28)
We obtain the third inequality by invoking Assumption 1. The fourth inequality arises from the fact that $F_{i,S}(w,\cdot)$ is $\rho$-strongly concave. The fifth inequality is obtained from the property of 1. The final inequality follows from the bound $\eta_{\lambda,t}\leq\frac{1}{L}$. By combining equations (LABEL:youngone) and (28) with equation (26), we arrive at the desired result. ∎
@@
Define $a_{t}=\mathbb{E}[R_{S}(w_{t})-R^{*}_{S}]$ and $b_{t}=\mathbb{E}\|\lambda_{t}-\lambda^{*}_{S}(w_{t})\|^{2}$. For any non-increasing sequence $\{\nu_{t}>0\}$ and any $\epsilon>0$, we have the following relations,
$a_{t+1}+\nu_{t+1}b_{t+1}\leq k_{1,t}a_{t}+k_{2,t}\nu_{t}b_{t}+c_{t},$ (29)
@@
$k_{1,t}$ $=1-\frac{\mu\eta_{w,t}}{\beta^{2}}$ (30)
$k_{2,t}$ $=\frac{EL^{2}\eta_{w,t}}{2\nu_{t}}+(1+\epsilon)(1-\rho\eta_{\lambda,t})$ (31)
$c_{t}$ $=-\frac{\mu}{\beta^{2}}\eta_{w,t}\Gamma+\frac{\eta_{w,t}}{2\beta^{2}}\kappa^{2}+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}E^{2}B_{w}^{2}$
$+\frac{\eta_{w,t}^{3}L^{2}E(E+1)}{4}(B_{w}^{2}+G^{2})$
$+\nu_{t}(1+\frac{1}{\epsilon})\frac{L^{2}E\eta_{w,t}^{2}}{\rho^{2}}(B_{w}^{2}+G^{2})+\nu_{t}(1+\epsilon)\eta_{\lambda,t}^{2}B_{\lambda}^{2}.$ (32)
@@
$k_{2,t}=\frac{EL^{2}\eta_{w,t}}{2\nu_{t}}+1-\frac{\rho\eta_{\lambda,t}}{2}.$
@@
$k_{1,t}$ $=1-\frac{\mu\eta_{w,t}}{\beta^{2}}$ (33)
@@
By setting $\eta_{\lambda,t}\geq\frac{2\mu}{\rho\beta^{2}}\eta_{w,t}$, we get
$a_{t+1}+\nu_{t+1}b_{t+1}\leq(1-\frac{\mu}{2\beta^{2}}\eta_{w,t})(a_{t}+\nu_{t}b_{t})-\frac{\mu}{\beta^{2}}\eta_{w,t}\Gamma+\frac{\eta_{w,t}}{2\beta^{2}}\kappa^{2}$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}E^{2}B_{w}^{2}+\frac{\eta_{w,t}^{3}L^{2}E(E+1)}{4}(B_{w}^{2}+G^{2})$
$+\nu_{t}(1+\frac{1}{\epsilon})\frac{L^{2}E\eta_{w,t}^{2}}{\rho^{2}}(B_{w}^{2}+G^{2})+\nu_{t}(1+\epsilon)\eta_{\lambda,t}^{2}B_{\lambda}^{2}.$ (35)
Choose $\eta_{w,t}=\frac{2\beta^{2}}{\mu t}$ and $\eta_{\lambda,t}=\min\{\frac{1}{L},\frac{4}{\rho t^{2/3}}\}$ and multiply both sides of (35) by $t$ to get
$t(a_{t+1}+\nu_{t+1}b_{t+1})$ $\leq(t-1)(a_{t}+\nu_{t}b_{t})+\frac{4\beta^{4}E^{2}B_{w}^{2}(L+\frac{L^{2}}{\rho})}{\mu^{2}t}$
$+\frac{L^{2}E(E+1)8\beta^{6}}{\mu^{3}t^{2}}(B_{w}^{2}+G^{2})$
$+\frac{2L^{4}E\beta^{6}}{\rho^{2}\mu^{3}t^{2/3}}+\frac{4\beta^{2}L^{2}}{\mu\rho^{2}t^{2/3}}+\frac{4\beta^{2}L^{2}}{\mu\rho^{2}(t^{2/3}-4)}$
$+\frac{\kappa^{2}}{\mu}-\frac{\Gamma}{2}.$ (36)
Applying the inequality inductively from $t=0$ to $T$ and divide both sides by $T$, we get
$\mathbb{E}[R_{S}(w_{T+1})-R^{*}_{S}]$ $\leq\frac{4\beta^{4}E^{2}B_{w}^{2}(L+\frac{L^{2}}{\rho})\ln T}{\mu^{2}T}$
$+\frac{L^{2}E(E+1)8\beta^{6}}{\mu^{3}T^{2}}(B_{w}^{2}+G^{2})+\frac{6L^{4}E\beta^{6}}{\rho^{2}\mu^{3}T^{2/3}}$
$+\frac{12\beta^{2}L^{2}}{\mu\rho^{2}T^{2/3}}+\frac{12\beta^{2}L^{2}}{\mu\rho^{2}(T^{2/3}-4)}$
$+\frac{\kappa^{2}}{\mu}-\frac{\Gamma}{2}.$ (37)
Appendix B Analysis of The Upper Bound of The Convergence Rate on a Simple Example
The theoretical result shows that on nonconvex setting, the algorithm could converge to a different (non-optimal) saddle point. This can be understood from a simple example consisting of two clients with the following local risk functions
$F_{1}(w,\lambda)$ $=(w-1)^{2}+2sin^{2}(w-2)-(\lambda-4)^{2}$
$F_{2}(w,\lambda)$ $=(w+1)^{2}+2.1sin^{2}(w+2)-(\lambda+4)^{2}.$ (38)
Also, define the global risk function as $F_{S}(w,\lambda)=0.5F_{1}(w,\lambda)+0.5F_{2}(w,\lambda)$. It can be shown that $F_{1}^{*}\approx 0.657645$, $F_{2}^{*}\approx 0.669149$, and $(w^{*},\lambda^{*})\approx(0.650928,0)$ with $F^{*}\approx-13.3915$. There exists another saddle point where it satisfies the second-derivative test at $(w,\lambda)\approx(-0.633757,0)$ with $F(w,\lambda)=-13.3552$. It can also be shown that by restricting the domain, $w\in[-3,3]$ and $\lambda\in[-5,5]$, we have $\mu=\frac{1}{8},\rho=1,L=18,G=11,\kappa^{2}=3.7^{2},\beta^{2}=1,\Gamma\approx-14.72$. This means that the nonvanishing term has a nonzero positive dependency with the heterogeneity parameters. If we run FFALM with $(w_{0},\lambda_{0})=(2,2)$ for 1 million communication rounds, we get $(w_{T},\lambda_{T})\approx(-0.6338911,0)$. This shows that the algorithm may converge to different saddle point locations.
Appendix C Additional Experiments on Different Levels of Heterogeneity